OpenStudy (anonymous):

find the dimensions of the largest rectangle that can be inscribe in the region bounded by f(x)=-x^2+4 and g(x)=x^2-4

OpenStudy (anonymous):

\[ Area(x)= 2 x ( 2(-x^2 +4)=16 x-4 x^3\\ Area'(x)=16 -12 x^2\\ Area'(x)= 0 \text{ for } x= \frac{2}{\sqrt{3}} \text { and } x=-\frac{2}{\sqrt{3}}\\ Area''( \frac{2}{\sqrt{3}})= -16 \sqrt{3} <0\\ \] The dimensions are;; \[ 2 \frac{2}{\sqrt{3}} = \frac{4}{\sqrt{3}}\\ 2 ( -\left(\frac{2}{\sqrt{3}})\right)^2 +4=\frac {16} 3 \]